# RDA-INR: Riemannian Diffeomorphic Autoencoding via Implicit Neural Representations

### v2

@@ 4 Riemannian Diffeomorphic Autoencoding via Implicit Neural Representations
Figure 3: Overview of our RDA-INR framework for resolution independent and physically-consistent statistical latent modeling and atlas building. Our method maps a latent vector $z_{i}$ to an INR representing a time-dependent velocity vector field $v_{\varphi}(\cdot,t,z_{i})$. This velocity vector field defines a flow of diffeomorphisms $\phi_{t}^{z_{i}}$ via an ODE. The diffeomorphisms $\phi_{1}^{z_{i}}$ create (reconstructed) objects $O_{i}$ on a subset of the Riemannian LDDMM manifold $\mathcal{M}$ by deforming a learned template $\mathcal{T}_{\theta}$. We obtain this deformation by parameterizing $\mathcal{T}_{\theta}$ with an INR $f_{\theta}$ and deforming $f_{\theta}$ via a group action: $I_{i}(\cdot,t):=(\phi_{t}^{z_{i}})^{-1}\cdot f_{\theta}$. As the template and the velocity vector fields are parameterized by INRs, $I_{i}$ is a 4D INR deforming the template at $t=0$ to a reconstruction at $t=1$. The goal is to learn the template $\mathcal{T}_{\theta}$ as Fréchet mean of the data and to learn the $I_{i}(\cdot,t)$ paths as geodesics on $\mathcal{M}$ between the template $\mathcal{T}_{\theta}$ and the data.
Our model depicted in Figure 3 can be obtained similarly to how we derived LDDMM PGA in Section 3.3. Specifically, we derive our model by starting at the LDDMM Fréchet mean problem [32, 82, 33, 34].
@@
$\min_{\{v_{i}(\cdot,t)\}_{i=1}^{N},\mathcal{T}}\quad$ $\frac{1}{N}\sum_{i=1}^{N}\left[\mathcal{D}\left(\phi_{1}^{i}\cdot\mathcal{T},O_{i}\right)+\sigma^{2}\int_{0}^{1}\lVert v_{i}(\cdot,t)\rVert_{V}^{2}\mathrm{d}t\right]$ (11)
@@
For LDDMM PGA, we required discretization of the template and initial velocity vector fields to solve the optimization problem in Equation (11). Furthermore, the discretization allows the restriction of the initial velocity vector fields to a linear subspace. Our approach circumvents the need for discretization by parameterizing the template and time-dependent velocity vector fields via INRs. Specifically, we parameterize $\phi_{1}^{i}$ by specifying the corresponding flow in reverse time. Defining $z_{i}\in\mathbb{R}^{d_{z}}$ as the latent code belonging to the object $O_{i}$ and $v_{\varphi}:\Omega\times[0,1]\times\mathbb{R}^{d_{z}}\rightarrow\mathbb{R}^{d}$ as an implicit neural representation, the reverse time flow is given by $\frac{\partial}{\partial t}\phi_{t}^{z_{i}}(x)=v_{\varphi}(\phi_{t}^{z_{i}}(x),t,z_{i})$. Jointly learning the weights $\varphi$ and the latent codes $z_{i}$ gives:
$\min_{\varphi,\{z_{i}\}_{i=1}^{N},\mathcal{T}}\quad$ $\frac{1}{N}\sum_{i=1}^{N}\left[\mathcal{D}((\phi_{1}^{z_{i}})^{-1}\cdot\mathcal{T},O_{i})+\sigma^{2}\int_{0}^{1}\lVert v_{\varphi}(\cdot,t,z_{i})\rVert_{V}^{2}\mathrm{d}t\right]$ (12)
@@
Finally, we represent the template $\mathcal{T}$ by an implicit neural representation to obtain infinite-resolution reconstructions. Specifically, when dealing with images we represent $\mathcal{T}$ as a neural network $f_{\theta}:\Omega\rightarrow\mathbb{R}^{n}$ and for shapes we represent it with the zero-level set of a neural network $f_{\theta}:\Omega\rightarrow\mathbb{R}$. In both cases, the group action on the implicit representation becomes:
$I(x,z,t):=((\phi_{t}^{z})^{-1}\cdot f_{\theta})(x):=f_{\theta}(\phi_{t}^{z}(x)),$ (13)
where $I(x,z,0)$ equals the implicit representation of the template and $I(x,z,1)$ is the reconstructed implicit representation of an object $O$. When dealing with shapes, we can obtain a mesh and point cloud for the reconstructed shape by applying marching cubes on $I(x,z,1)$. Another strategy is to apply marching cubes to the template INR $f_{\theta}$ to obtain a template mesh $\mathcal{M}_{\mathcal{T}}$ and then retrieve the reconstructed mesh $\mathcal{M}_{z}$ by $(\phi_{1}^{z})^{-1}(\mathcal{M}_{\mathcal{T}})$.
Note that the group action in Equation (13) is a group action on the implicit representation. When dealing with shapes, the group action in Equation (13) can be associated to a group action on the template shape $\mathcal{T}_{\theta}:=\{x\mid f_{\theta}(x)=0\}$, namely $\phi\cdot\mathcal{T}_{\theta}:=\phi(\mathcal{T}_{\theta})$. In the case of images, $\mathcal{T}_{\theta}=f_{\theta}$ and we can use the group action in Equation (13). Incorporating the template $\mathcal{T}_{\theta}$ and the associated group action into Equation (12) yields the final optimization problem:
$\min_{\theta,\varphi,\{z_{i}\}_{i=1}^{N}}\quad$ $\frac{1}{N}\sum_{i=1}^{N}\left[\mathcal{D}((\phi_{1}^{z_{i}})^{-1}\cdot\mathcal{T}_{\theta},O_{i})+\sigma^{2}\int_{0}^{1}\lVert v_{\varphi}(\cdot,t,z_{i})\rVert_{V}^{2}\mathrm{d}t\right]$ (14)
@@
Comparing Equation (14) with the LDDMM PGA optimization problem in Equation (10), we see several similarities and differences. In both problems, we optimize over latent vectors $z_{i}$ and weights to define diffeomorphisms. However, in LDDMM PGA, the weights construct a linear relationship between the latent vectors and discrete initial velocity vector fields, while the weights in our model parameterize neural networks. This parameterization yields a nonlinear relationship between latent vectors and the time-dependent velocity vector fields. Moreover, the velocity vector fields and the template are not discretized in our model. Hence, our model still views the velocity vector fields and the template as elements of infinite-dimensional spaces, making the model resolution-independent. Finally, due to the nonlinear relationship between latent vectors and velocity vector fields, and the infinite-dimensional nature of the vector fields, it is not possible to enforce the orthogonality properties of LDDMM PGA using an orthogonal weight matrix $W$ as shown in Equation (10).
Up until now, we kept the data fidelity term $\mathcal{D}$ general such that the optimization problem in Equation (14) applies to both images and shapes. In the case of images $I_{1}$ and $I_{2}$, $\mathcal{D}(I_{1},I_{2})=\lVert I_{1}-I_{2}\rVert_{2}^{2}$ is an appropriate data fidelity. When considering point cloud or mesh data without point correspondences, it is not immediately obvious what data fidelity $\mathcal{D}$ to use in combination with our INR representation of the template. As LDDMM PGA is focused on image data, our numerical experiments focus on the novel point cloud and mesh data. We introduce the used data fidelity term for this data in Section 4.1. After introducing the data fidelity term, we introduce the chosen $\lVert\cdot\rVert_{V}$ in Section 4.2. Subsequently, in Section 4.3, we introduce the neural network parameterization of $v_{\varphi}$ and we treat how we numerically deform the template to obtain a reconstruction. Finally, we treat how we can obtain latent codes for unseen instances in Section 4.4.
@@ 4.1 Data fidelity term for shape data
When dealing with shapes in the form of point cloud or mesh data, recent resolution-independent methods employ a deformable template for joint shape encoding and registration [39, 21] and use $\mathcal{D}(I_{1},I_{2})=\lVert I_{1}-I_{2}\rVert_{L_{1}(\Omega)}$ with $I_{i}$ a ground truth signed distance representation of the point cloud or mesh. However, there are two issues with such a $\mathcal{D}$: (i) signed distance functions of diffeomorphic shapes are not always diffeomorphic, and (ii) if they are diffeomorphic, they influence the diffeomorphism that we find.
To illustrate the first issue, we take two circles of radii $0.75$ and $0.1$, respectively. Moreover, assume they are represented by their SDF, as presented in Equation (1). These SDFs are shown in Figure 4. We note that $\min_{x\in\Omega}\text{SDF}_{\mathcal{S}_{0}}(x)<\min_{x\in\Omega}\text{SDF}_{\mathcal{S}_{1}}(x)$ as dark blue is present in the figure of the circle with radius $0.75$ but not in the other figure. Consequently, there does not exist a diffeomorphism that matches the two implicit representations (i.e., SDF images). However, in Equation (14) we want to diffeomorphically match both to the same template implicit function. This matching is not possible as the two SDFs can not be diffeomorphically registered.
@@
To overcome these issues, we use the approach used by Sitzmann et al. [75], Deng et al. [38], and Gropp et al. [83] for learning an SDF from point cloud and mesh data. This approach uses the point cloud representation for training shapes $\{\mathcal{S}_{i}\}_{i=1}^{N}$ and uses the following data fidelity term $\mathcal{D}$:
$\mathcal{D}((\phi_{1}^{z})^{-1}\cdot\mathcal{T}_{\theta},\mathcal{S}_{i}):=\operatorname*{\mathop{\mathbb{E}}}_{x\in\mathcal{S}_{i}}\left[\lvert I(x,z,1)\rvert+\tau\left(1-F_{\text{cos}}(\nabla_{x}I(x,z,1),n_{i}(x))\right)\right],$
with $I(x,z,t)$ given in Equation (13), $\tau\in\mathbb{R}$, $n_{i}(x)$ the unit normal of the shape $\mathcal{S}_{i}$ at $x\in\mathcal{S}_{i}$, and $F_{\text{cos}}$ the cosine similarity:
$F_{\text{cos}}(a,b):=\frac{\langle a,b\rangle_{2}}{\max(\lVert a\rVert_{2},\lVert b\rVert_{2},10^{-8})},\quad a,b\in\mathbb{R}^{d}.$
This data-fitting term solves the previously mentioned issues since it only focuses on matching points on the training shapes with points on the template shape. In other words, all points on the training shapes are matched to the zero level set of the INR $f_{\theta}$. However, using only this loss function encourages $f_{\theta}=0$ as this ensures that all points are matched to the zero level set of $f_{\theta}$. To avoid this trivial template shape, we constrain the learned template in optimization problem (14) via additional loss terms:
$\min_{\theta,\varphi,\{z_{i}\}_{i=1}^{N}}\quad$ $\frac{1}{N}\sum_{i=1}^{N}\left[\operatorname*{\mathop{\mathbb{E}}}_{x\in\mathcal{S}_{i}}\left[\lvert I(x,z_{i},1)\rvert+\tau\left(1-F_{\text{cos}}(\nabla_{x}I(x,z_{i},1),n_{i}(x))\right)\right]\right.$ (15)
$\left.\qquad\quad+\beta\operatorname*{\mathop{\mathbb{E}}}_{x\in\Omega\setminus\mathcal{S}_{i}}\left[\exp{(-\alpha\lvert I(x,z_{i},1)\rvert)}\right]+\sigma^{2}\int_{0}^{1}\lVert v_{\varphi}(\cdot,t,z_{i})\rVert_{V}^{2}\mathrm{d}t\right]$
$+\lambda\operatorname*{\mathop{\mathbb{E}}}_{x\in\Omega}\left[\lvert\lVert\nabla_{x}f_{\theta}(x)\rVert_{2}-1\rvert\right]$
@@
Here the penalty corresponding to $\lambda\in\mathbb{R}$ is an eikonal penalty, regularizing the template shape to be a signed distance function as in Equation (1) and ensuring that $f_{\theta}\neq 0$. Furthermore, the loss function involving the $\beta$ parameter ensures that only points on $\mathcal{S}_{i}$ are matched to the zero level set of $f_{\theta}$. Both terms are used by Sitzmann et al. [75], Deng et al. [38], and Gropp et al. [83] for learning an SDF from point cloud and mesh data.
Remark 1.
Another approach to solve the issues is using the occupancy function in Equation (2) as the implicit representation for the data and the template shape. Earlier works using occupancy functions as implicit representations are Mescheder et al. [76] and Niemeyer et al. [68]. The reason occupancy functions solve the issues is that we only match points inside (outside) shape $\mathcal{S}_{0}$ to points inside (outside) shape $\mathcal{S}_{1}$ and do not take into account information like the SDF value. As the occupancy function and the occupancy values resemble probabilities, we can use the binary cross entropy loss as data fidelity term $\mathcal{D}$ in optimization problem (14) [76, 68]. However, as shown in Appendix B, our strategy that uses point clouds as data representation allows for higher-quality reconstructions. Consequently, we use the point cloud data representation instead of the occupancy value data representation. Another reason for using point cloud data is related to the shape representation used by diffeomorphic registration methods. Specifically, these methods use either (i) meshes and point clouds or (ii) occupancy functions/segmentation masks. Our approach can deal with both data types as (i) our method deals with point cloud data via an image-based technique, and (ii) we can deal with segmentation masks via occupancy functions. We are showcasing the point cloud approach as, to the best of our knowledge, it is the first approach for dealing with point clouds using image-based representations.
4.2 Choice of velocity field regularization term
For the norm $\lVert\cdot\rVert_{V}$ in optimization problems (14) and (15) we choose an isometric (rigid) deformation prior on the velocity vector fields by combining the Killing energy [84, 85] with an $L_{2}(\Omega)$ penalty:
$\lVert\nu\rVert_{V}^{2}:=\int_{\Omega}\lVert(J\nu)+(J\nu)^{T}\rVert_{F}^{2}+\eta\lVert\nu\rVert^{2}_{2}\mathrm{d}x,$ (16)
where $\nu$ is a velocity vector field, $J\nu$ the Jacobian of $\nu$ with respect to $x$, and $\eta\in\mathbb{R}$. If $\eta$ and the functional in Equation (16) are small, then we expect the template to deform almost isometrically to the reconstructed objects via $I(x,z,t)$ in Equation (13). This property is interesting as it allows us to jointly perform affine registration and LDDMM registration. Moreover, as shown in Appendix C, under certain conditions the Killing energy can be viewed as an extension of a standard norm used in LDDMM.
4.3 Solving the ordinary differential equation
$v_{\varphi}(x,t,z)=\sum_{k=1}^{K}\chi_{[\frac{k-1}{K},\frac{k}{K})}(t)\cdot v_{\varphi_{k}}(x,z).$ (17)

### v3

@@ 4 Riemannian Diffeomorphic Autoencoding via Implicit Neural Representations
Figure 3: RDA-INR framework for resolution-independent and physically-consistent statistical latent modeling and atlas building. Our method maps a latent vector $z_{i}$ to an INR representing a time-dependent vector field $v_{\varphi}(\cdot,t,z_{i})$. This vector field defines a flow of diffeomorphisms $\phi_{t}^{z_{i}}$ via an ODE. The diffeomorphisms $\phi_{1}^{z_{i}}$ create (reconstructed) objects $O_{i}$ on a subset of the Riemannian LDDMM manifold $\mathcal{M}$ by deforming a learned template $\mathcal{T}_{\theta}$. We obtain this deformation by parameterizing $\mathcal{T}_{\theta}$ with an INR $f_{\theta}$ and deforming $f_{\theta}$ via a group action: $I_{i}(\cdot,t):=(\phi_{t}^{z_{i}})^{-1}\cdot f_{\theta}$. As the template and the vector fields are parameterized by INRs, $I_{i}$ is a 4D INR deforming the template at $t=0$ to a reconstruction at $t=1$. The goal is to learn the template $\mathcal{T}_{\theta}$ as Fréchet mean of the data and to learn the $I_{i}(\cdot,t)$ paths as geodesics on $\mathcal{M}$ between the template $\mathcal{T}_{\theta}$ and the data.
Our model depicted in Figure 3 can be obtained similarly to how we derived LDDMM PGA in Section 3.3. Specifically, we derive our model by starting with the LDDMM Fréchet mean problem [40, 10, 46, 47].
@@
(9) $\min_{\{v_{i}(\cdot,t)\}_{i=1}^{N},\mathcal{T}}$ $\frac{1}{N}\sum_{i=1}^{N}\left[\mathcal{D}\left(\phi_{1}^{i}\cdot\mathcal{T},O_{i}\right)+\sigma^{2}\int_{0}^{1}\left\lVert v_{i}(\cdot,t)\right\rVert_{V}^{2}\mathrm{d}t\right]$
@@
For LDDMM PGA, we required a grid-based or spatial discretization of the template and initial velocity vector fields to solve the optimization problem in Equation (9). Furthermore, the chosen discretization allows the restriction of the initial velocity vector fields to a linear subspace. Our approach circumvents the need for grid-based or spatial discretizations by parameterizing the template and time-dependent velocity vector fields via INRs. Specifically, we parameterize $\phi_{1}^{i}$ by specifying the corresponding flow in reverse time. Defining $z_{i}\in\mathbb{R}^{d_{z}}$ as the latent code belonging to the object $O_{i}$ and $v_{\varphi}\colon\Omega\times[0,1]\times\mathbb{R}^{d_{z}}\rightarrow\mathbb{R}^{d}$ as an implicit neural representation, the reverse time flow is given by $\frac{\partial}{\partial t}\phi_{t}^{z_{i}}(x)=v_{\varphi}(\phi_{t}^{z_{i}}(x),t,z_{i})$. Jointly learning the weights $\varphi$ and the latent codes $z_{i}$ gives:
(10) $\min_{\varphi,\{z_{i}\}_{i=1}^{N},\mathcal{T}}$ $\frac{1}{N}\sum_{i=1}^{N}\left[\mathcal{D}((\phi_{1}^{z_{i}})^{-1}\cdot\mathcal{T},O_{i})+\sigma^{2}\int_{0}^{1}\left\lVert v_{\varphi}(\cdot,t,z_{i})\right\rVert_{V}^{2}\mathrm{d}t\right]$
@@
Finally, we represent the template $\mathcal{T}$ by an implicit neural representation $f_{\theta}\colon\Omega\rightarrow\mathbb{R}$ to obtain a resolution-independent method. The group action on the implicit neural representation is:
(11) $I(x,z,t):=((\phi_{t}^{z})^{-1}\cdot f_{\theta})(x):=f_{\theta}(\phi_{t}^{z}(x)),$
where $I(x,z,0)$ equals the implicit representation of the template and $I(x,z,1)$ is the reconstructed implicit representation of an object $O$. This group action motivates parameterizing the flow in reverse time as we need to evaluate $f_{\theta}$ on the reverse-time flow. Alternatively, the group action can be associated with a group action on the set-based definition of the template shape $\mathcal{T}_{\theta}:=\{x\mid f_{\theta}(x)=0\}$, namely $\phi\cdot\mathcal{T}_{\theta}:=\phi(\mathcal{T}_{\theta})$. Both group actions yield a strategy to obtain a reconstructed mesh or point cloud for the reconstructed shape: apply marching cubes on $I(x,z,1)$ or use $(\phi_{1}^{z})^{-1}\cdot\mathcal{T}_{\theta}=(\phi_{1}^{z})^{-1}(\mathcal{T}_{\theta})$. Incorporating the template $\mathcal{T}_{\theta}$ and the associated group action into Equation (10) and assuming some $\mathcal{D}$ for fitting the deformed template or its implicit representation to the data, yields the final optimization problem:
(12) $\min_{\theta,\varphi,\{z_{i}\}_{i=1}^{N}}$ $\frac{1}{N}\sum_{i=1}^{N}\left[\mathcal{D}((\phi_{1}^{z_{i}})^{-1}\cdot\mathcal{T}_{\theta},O_{i})+\sigma^{2}\int_{0}^{1}\left\lVert v_{\varphi}(\cdot,t,z_{i})\right\rVert_{V}^{2}\mathrm{d}t\right]$
@@
Comparing Equation (12) with the LDDMM PGA optimization problem in Equation (8), we see several similarities and differences. In both problems, we optimize over latent vectors $z_{i}$ and weights to define diffeomorphisms. However, in LDDMM PGA, the weights construct a linear relationship between the latent vectors and discrete initial velocity vector fields, while the weights in our model parameterize neural networks. This parameterization yields a nonlinear relationship between latent vectors and the time-dependent velocity vector fields. Compared to the linear relationship, the nonlinear relationship makes the latent space more difficult to navigate as it is a-priori unclear how a change in latent space affects the velocity vector fields. Moreover, our model does not use a grid-based or spatial discretization for the velocity vector fields and the template. Instead, our model parameterizes them via INRs, making the model resolution independent. By optimizing the weights of the INRs, the model searches over finite-dimensional nonlinear submanifolds of full infinite-dimensional spaces. Specifically, within these submanifolds, our model searches for the best template and nonlinear submanifold of time-dependent velocity vector fields. Finally, due to the nonlinear relationship between latent vectors and velocity vector fields, and the INR representation of the vector fields, it is not possible to enforce the orthogonality properties of LDDMM PGA using an orthogonal weight matrix $W$ as shown in Equation (8).
Up until now, we kept the data fidelity term $\mathcal{D}$ general such that the optimization problem in Equation (12) is as close as possible to the general LDDMM PGA model. For the point cloud or mesh data that we consider, it is not immediately obvious what data fidelity $\mathcal{D}$ to use in combination with our INR representation of the template. We introduce the used data fidelity term for this data in Section 4.1. Subsequently, Section 4.2 introduces the velocity field parameterization and the chosen $\left\lVert\cdot\right\rVert_{V}$. Finally, we treat how we obtain latent codes for unseen instances in Section 4.3. For details about the neural network architectures of $f_{\theta}$ and $v_{\varphi}$ and how we numerically deform the template to obtain a reconstruction, see Appendix D.
@@ 4.1 Data fidelity term for shape data
When dealing with point cloud or mesh data, recent resolution-independent methods employ a deformable template for joint shape encoding and registration [85, 63] and use $\mathcal{D}(I_{1},I_{2})=\left\lVert I_{1}-I_{2}\right\rVert_{L_{1}(\Omega)}$ with $I_{i}$ a ground truth signed distance representation of the point cloud or mesh. However, there are three issues with such a $\mathcal{D}$: (i) signed distance functions of diffeomorphic shapes are not always diffeomorphic, (ii) if they are diffeomorphic, they influence the diffeomorphism that we find, and (iii) signed distance functions are not always available for point clouds or are non-trivial to obtain.
To illustrate the first issue, take two circles of radii $0.75$ and $0.1$, respectively. Moreover, represent them by their SDF, as presented in Equation (1). These SDFs are shown in Figure 4. We have $\min_{x\in\Omega}\text{SDF}_{\mathcal{S}_{0}}(x)<\min_{x\in\Omega}\text{SDF}_{\mathcal{S}_{1}}(x)$ as dark blue is present in the figure of the circle with radius $0.75$ but not in the other figure. Consequently, no diffeomorphism matching the two implicit representations (i.e., SDF images) exists, which contradicts Equation (12) where both SDFs need to be diffeomorphically matched to the same template implicit function.
@@
To overcome these issues, we use the approach used by Sitzmann et al. [61] and Gropp et al. [31] for learning an SDF from point cloud and mesh data. This approach uses the point cloud representation for training shapes $\{\mathcal{S}_{i}\}_{i=1}^{N}$ and uses the following data fidelity term $\mathcal{D}$:
(13) $\mathcal{D}((\phi_{1}^{z})^{-1}\cdot\mathcal{T}_{\theta},\mathcal{S}_{i}):=\operatorname*{\mathop{\mathbb{E}}}_{x\in\mathcal{S}_{i}}\left[\lvert I(x,z,1)\rvert\right.$ $\left.+\tau\left(1-F_{\text{cos}}(\nabla_{x}I(x,z,1),n_{i}(x))\right)\right]$
$+\beta\operatorname*{\mathop{\mathbb{E}}}_{x\in\Omega\setminus\mathcal{S}_{i}}\left[\exp{(-\alpha\lvert I(x,z_{i},1)\rvert)}\right],$
with $I(x,z,t)$ given in Equation (11), $\tau,\beta,\alpha\in\mathbb{R}$, $n_{i}(x)$ the unit normal of the shape $\mathcal{S}_{i}$ at $x\in\mathcal{S}_{i}$, and $F_{\text{cos}}$ the cosine similarity:
$F_{\text{cos}}(a,b):=\frac{\langle a,b\rangle_{2}}{\max(\left\lVert a\right\rVert_{2}\cdot\left\lVert b\right\rVert_{2},10^{-8})},\quad a,b\in\mathbb{R}^{d}.$
This data-fitting term solves the previously mentioned issues since the first expectation only focuses on matching points on the training shapes with points on the template shape. In other words, all points on the training shapes are matched to the zero level set of the INR $f_{\theta}$. However, using only the first expectation encourages $f_{\theta}=0$ as this ensures all points are matched to the zero level set of $f_{\theta}$. To avoid this trivial template, the loss term involving the $\beta$ parameter ensures that only points on $\mathcal{S}_{i}$ are matched to the zero level set of $f_{\theta}$. In addition, to enforce additional structure on the learned template INR and to further ensure $f_{\theta}\neq 0$, we add an eikonal penalty [61, 25, 31] with regularization constant $\lambda\in\mathbb{R}$ to optimization problem (12), which regularizes the template INR to be a signed distance function as in Equation (1). Finally, we add an isotropic Gaussian prior with regularization constant $\gamma$ on the latent codes. Combining all loss functions results in the final optimization problem:
(14) $\min_{\theta,\varphi,\{z_{i}\}_{i=1}^{N}}$ $\frac{1}{N}\sum_{i=1}^{N}\left[\operatorname*{\mathop{\mathbb{E}}}_{x\in\mathcal{S}_{i}}\left[\lvert I(x,z_{i},1)\rvert+\tau\left(1-F_{\text{cos}}(\nabla_{x}I(x,z_{i},1),n_{i}(x))\right)\right]\right.$
$\left.\qquad\quad+\beta\operatorname*{\mathop{\mathbb{E}}}_{x\in\Omega\setminus\mathcal{S}_{i}}\left[\exp{(-\alpha\lvert I(x,z_{i},1)\rvert)}\right]+\sigma^{2}\int_{0}^{1}\left\lVert v_{\varphi}(\cdot,t,z_{i})\right\rVert_{V}^{2}\mathrm{d}t+\gamma\left\lVert z_{i}\right\rVert_{2}^{2}\right]$
$+\lambda\operatorname*{\mathop{\mathbb{E}}}_{x\in\Omega}\left[\lvert\left\lVert\nabla_{x}f_{\theta}(x)\right\rVert_{2}-1\rvert\right]$
@@
Remark 4.1.
Another approach to solve the issues is using the occupancy function in Equation (2) as the implicit representation for the data and the template shape. Earlier works using occupancy functions as implicit representations are Mescheder et al. [48] and Niemeyer et al. [53]. The reason occupancy functions solve the issues is that we only match points inside (outside) shape $\mathcal{S}_{0}$ to points inside (outside) shape $\mathcal{S}_{1}$ and do not take into account information like the SDF value. As the occupancy function and the occupancy values resemble probabilities, we can use the binary cross entropy loss as data fidelity term $\mathcal{D}$ in optimization problem (12) [48, 53]. However, as shown in Appendix B.1, our strategy that uses point clouds as data representation allows for higher-quality reconstructions. This result further motivates the use of the point cloud representation.
Remark 4.2.
Although we specialize our model to point cloud and mesh data representing shapes, our model can easily be extended to images that do not necessarily represent shapes. Specifically, in Equation (12), we can choose $\mathcal{D}(I_{1},I_{2})$ as any data fitting attachment for images and use $\mathcal{T}_{\theta}=f_{\theta}$ with $\phi\cdot\mathcal{T}_{\theta}=\mathcal{T}_{\theta}\circ\phi^{-1}$.
4.2 Choice of velocity field parameterization and velocity field regularization term
As parameterization of $v_{\varphi}$, we choose a quasi-time varying velocity vector field defined by several INRs representing stationary velocity vector fields. For more details, we refer to Appendix D. For the norm $\left\lVert\cdot\right\rVert_{V}$ in optimization problems (12) and (14) we choose an isometric (rigid) deformation prior on the velocity vector fields by combining the Killing energy [62, 64] with an $L_{2}(\Omega)$ penalty:
(15) $\left\lVert\nu\right\rVert_{V}^{2}:=\int_{\Omega}\lVert(J\nu)+(J\nu)^{T}\rVert_{F}^{2}+\eta\left\lVert\nu\right\rVert^{2}_{2}\mathrm{d}x,$
where $\nu$ is a velocity vector field, $J\nu$ the Jacobian of $\nu$ with respect to $x$, and $\eta\in\mathbb{R}$. If $\eta$ and the functional in Equation (15) are small, then we expect the template to deform almost isometrically to the reconstructed objects via $I(x,z,t)$ in Equation (11). In the numerical experiments, we use this property and a dataset with a rigidity prior to highlight the importance of a data-consistent regularizer. Furthermore, this property is interesting as it allows us to jointly perform affine registration and LDDMM registration. The regularizer and its property are novel in the LDDMM literature. In Appendix C we showcase that our regularizer provides a novel motivation and special case of a lesser-used norm in the LDDMM literature. We also show that our regularizer adds a rigidity prior to a more standard Sobolev LDDMM norm.